OpenStudy (anonymous):

Factor: 125x^3+27y^3

OpenStudy (anonymous):

125x^3 = (5x)^3 27x^3 = (3x)^3 \(a^3+b^3 = (a+b)(a^2 - ab+b^2)\) Can you use this identity to do it?

OpenStudy (anonymous):

Just note that 125x^3 = (5x)^3

OpenStudy (anonymous):

and 27y^3 = (3y)^3

OpenStudy (anonymous):

I never used this formula in that sense before... with that I got 5^3x^3 + 3^3y^3, how would I plug that in with a and b?

OpenStudy (anonymous):

no,not like that

OpenStudy (anonymous):

(5x + 3y)(___ - __*__ + ____)

OpenStudy (anonymous):

Hey Ray - do you see the light, or to explain more ?

OpenStudy (anonymous):

Yeah, just give me a second to work it out and see if I come up with the right answer.

OpenStudy (anonymous):

All right Ray

OpenStudy (anonymous):

Okay, I got it, (5x+3y)(25x^2-15xy+9y^2)

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

OO kay , by the way what is your actual name (reply in message )

OpenStudy (anonymous):

I mean just the given name - not family